OpenStudy (anonymous):

please help me find the vertex and intercepts fory =2x^2-3x+3

OpenStudy (dumbcow):

x = -b/2a x = 3/4 plug it in to get y_coordinate vertex: (3/4, f(3/4)) For intercepts plug in 0 for x,y respectively y_int = (0,3) x_int --> 2x^2 -3x +3 = 0 use quadratic formula x = 3 +-sqrt(9-4(2)(3)) / 4 you get neg inside sqrt, no real solutions

OpenStudy (anonymous):

thank you...I didnt know how to begin it ,,,so the vertex is just (3/4)?

OpenStudy (anonymous):

is it not -b/2a? or is just b/2a ok?

OpenStudy (dumbcow):

no it has to be -b/2a and for the vertex (x,y) there is x_coord and y_coord just like any other point the formula -b/2a is a shortcut to find the x_coord, but you have to plug it in to the function to get the y_coord you can also use completing the square to obtain the vertex form, but many students do not like this method hence the shortcut